# Sample then Identify: A General Framework for Risk Control and Assessment in Multimodal Large Language Models

### v2

@@ 1 Introduction
In recent years, there has been swift progress in the development of Large Language Models (LLMs) and Multimodal Large Language Models (MLLMs) (Zhao et al., 2023; Wu et al., 2023a). MLLMs extend the capabilities of LLMs by integrating and processing information from multiple modalities, including text, vision, and audio (Bai et al., 2024; Zhu et al., 2024). However, MLLMs are proven to exhibit significant drawbacks in trustworthiness, such as hallucination (Rawte et al., 2023; Zhang et al., 2024; Wang et al., 2024a), which results in non-factual information (Ji et al., 2023) and biased generations (Zou et al., 2023). These issues have elicited increasing societal concerns regarding the reliable deployment of foundation models in consumer-facing applications (Qian et al., 2024).
Uncertainty estimation (Ye et al., 2024) provides valuable insights into the trustworthiness of model generations. Prior work performs probabilistic modeling (Li et al., 2024), develops entropy-based measures (Wang et al., 2024b; Duan et al., 2024; Farquhar et al., 2024), and generates multiple responses to analyze the output space (Lin et al., 2024; Wang et al., 2024c). However, these approaches cannot provide guarantees of the error rate, and forcing a measure to predict a single data point is overly restrictive (Cresswell et al., 2024). Split Conformal Prediction (SCP) (Papadopoulos et al., 2002; Angelopoulos and Bates, 2021; Campos et al., 2024) is a promising candidate to tackle these challenges above, which constructs statistically rigorous prediction sets with the hallmark of risk control (Angelopoulos et al., 2024; Farinhas et al., 2024). Furthermore, SCP allows the calibrated set size to vary based on the model’s uncertainty regarding a particular input.
SCP has been successfully applied in language modeling (Campos et al., 2024). However, existing methods either (1) modify the original model outputs to ensure factuality with desired probability (Mohri and Hashimoto, 2024; Cherian et al., 2024), potentially leading to uninformative and vague generations, (2) rely on internal token sequence logits (Quach et al., 2024), or (3) are restricted to multiple-choice settings (Kumar et al., 2023; Ye et al., 2024; Kostumov et al., 2024). Adapting SCP for proprietary MLLMs in practical open-ended video question-answering (VideoQA) applications presents several challenges: (a) Adaptability. Unlike closed-ended tasks equipped with fixed options covering the correct answers, the output space for each VideoQA task is unbounded, and there may not be an acceptable response within the sampled generations; (b) Reliability. Logits or verbalized confidence levels are often miscalibrated, leading to biased prediction sets; (c) Flexibility. For some API-only MLLMs, users lack access to internal model information like logits.
@@
Once the risk level for acceptable responses being sampled is specified, we evaluate the reliability of each response. Inspired by self-consistency theory (Wang et al., 2022; Li et al., 2022a), we propose using the frequency of each response as a proxy for confidence [^1] Response frequency positively correlates with average real generative probability. As the frequency increases, there is a corresponding growing level of confidence in the model’s responses (Su et al., 2024)., compensating for the limitations of internal logits information and verbalized confidence scores, and define the nonconformity score of each calibration data as one minus the frequency score of any one acceptable response in the candidate set. Then, we specify another risk level $\beta$ and calculate the approximate $1-\beta$ quantile of all nonconformity scores, denoted as $\hat{s}$. The quantile serves as a threshold to identify high-quality responses within the candidate set of each test data point. Finally, we guarantee that the miscoverage rate of acceptable responses in the calibrated prediction sets is statistically controlled by two risk levels, $\alpha$ and $\beta$ (i.e., $\leq\alpha+\beta-\alpha\beta$), while the set size reflects the model uncertainty to each query.
We evaluate our framework on two closed-ended VideoQA datasets (e.g., Video-MME (Fu et al., 2024) and NExT-QA (Xiao et al., 2021)), and two open-ended VideoQA datasets (e.g., MUSIC-AVQA (Li et al., 2022b) and MSVD (Chen and Dolan, 2011)), utilizing five open-source MLLMs(e.g., VideoLLaMA (Zhang et al., 2023), PandaGPT (Su et al., 2023), and NExT-GPT (Wu et al., 2023b)), and three closed-source MLLMs(e.g., Gemini (Reid et al., 2024) and GPT-4o mini (OpenAI, 2023)). Experimental results demonstrate that TRON guarantees the error rates in both steps across various user-specified risk levels. Additionally, the deduplicated average set size provides stable uncertainty estimation of MLLMs across various risk levels in open-ended VideoQA tasks, complementing the accuracy metric for more comprehensive evaluations. Furthermore, we investigate other black-box measurements in the development of the nonconformity score, resulting in more efficient predictions and consistently rigorous guarantees. To summarize, our contributions are the following:
@@ 2 Related Work
Our work extends the variants of the SCP framework for correctness (coverage) guarantees and risk control in language modeling. Existing methods either modify model generations and achieve conformal factuality guarantees (Mohri and Hashimoto, 2024; Cherian et al., 2024) or guarantee the ground-truth coverage in Multiple-Choice Question-Answering (MCQA) tasks (Kumar et al., 2023; Ye et al., 2024; Kostumov et al., 2024). The previous approaches may result in vague and uninformative responses, constrained by the correctness of the original generation, while the latter frameworks are limited to closed-ended settings. In open-ended scenarios, (Quach et al., 2024) develops a stopping rule with slightly relaxed coverage criteria utilizing the internal logits information, restricting the applicability to API-only models, whereas (Su et al., 2024) leverages black-box uncertainty but fails to align the nonconformity score with correctness, leading to theoretically loose guarantees. Our approach addresses these gaps by introducing a novel two-step framework that first derives the minimum number of response samples to approximate a “closed-ended” condition, where there is at least one acceptable option provided, followed by identifying reliable responses based on self-consistency theory, which features (1) rigor, controlled by two user-specified risk levels, $\alpha$ and $\beta$, and (2) versatility, applicable to any MLLM that supports sampling in both closed-ended and open-ended settings.
@@
Recent work performs probabilistic modeling (Li et al., 2024), develops entropy-based uncertainty measures (Wang et al., 2024b; Duan et al., 2024; Farquhar et al., 2024), and generate multiple responses to analyze the output space (Lin et al., 2024; Wang et al., 2024c). However, these approaches are heuristic, and forcing a measure to predict a single data point is overly restrictive (Cresswell et al., 2024). (Ye et al., 2024) and (Kostumov et al., 2024) apply SCP to MCQA tasks, creating conformal sets based on a measure of uncertainty (risk), where the set size represents the confidence of the foundational model regarding a particular input, with larger sets indicating higher risk. In open-ended scenarios, there may be redundant responses within the prediction sets due to the randomness of sampling (Quach et al., 2024). Our work explores the performance of set size on risk assessment before and after deduplication and observes that the original set size becomes unstable in uncertainty estimation across different MLLMs as the error rate changes, while the calibrated set size reliably evaluates MLLMs across various risk levels.
@@ 3.1 Preliminaries
Following standard SCP for risk control (Angelopoulos and Bates, 2021; Angelopoulos et al., 2024), we assume a calibration dataset ${\mathcal{D}}_{cal}=\left\{\left(X_{i},Y_{i}\right)\right\}_{i=1}^{N}$ of $N$ query-answer pairs and an independent and identically distributed (i.i.d.) [^2] Split conformal prediction requires that the test data points are drawn from the same distribution as the calibration data points. test point $\left(x_{test},y_{test}\right)$. Our goal of risk control is to construct prediction sets that are calibrated to bound the correctness miscoverage rate
@@ 3.2 Conformal Risk Control
Inspired by self-consistency theory (Wang et al., 2022; Li et al., 2022a; Wang et al., 2024c), which states that a repetitively sampled response is viewed as a form of consistency linked to higher confidence in the response, we perform a simple semantic clustering process on the candidate set and obtain the frequency of each response, denoted as $F\left(\hat{y}_{m}^{(i)}\right)$. For a detailed description of semantic clustering, refer to Appendix C.1. Then, we define the nonconformity score of the $i$-th calibration data point as one minus the frequency of the acceptable response, denoted as $\hat{y}_{ref}^{(i)}$, within the candidate set, which is semantically equivalent to the label $Y_{i}$ (i.e., $\hat{y}_{ref}^{(i)}\Leftrightarrow Y_{i}$). $A\Leftrightarrow B$ represents a bidirectional entailment (MacCartney and Manning, 2015; Mohri and Hashimoto, 2024; Farquhar et al., 2024) between A and B (i.e., equivalence). The specific evaluation for bidirectional entailment is provided in Appendix C.1. In this case, the nonconformity score is formulated as
@@ 4.1 Experimental Settings
We employ five open-source MLLMs, including VideoLLaMA-7B (Zhang et al., 2023), VideoLLaMA-13B, PandaGPT-7B (Su et al., 2023), PandaGPT-13B, and NExT-GPT (Wu et al., 2023b), and three closed-source MLLMs, including Gemini-1.5-Flash (Reid et al., 2024), Gemini-1.5-Pro, and GPT-4o mini (OpenAI, 2023). Note that we utilize all five open-source MLLMs as if they are API-only MLLMs, i.e., it assumes no access to any internal information of MLLMs.
@@
We consider two multiple-choice VideoQA datasets: Video-MME (VMME) (Fu et al., 2024) and NExT-QA (NEXT) (Xiao et al., 2021), where we only utilize the multiple-choice section of NEXT, and two open-ended VideoQA datasets: MUSIC-AVQA (MUSC) (Li et al., 2022b) and MSVD (Chen and Dolan, 2011). More details of the dataset utilization can be found in Appendix C.
@@
We utilize the Empirical Error Rate (EER) to assess whether we control the risk of miscoverage by the two user-specified risk levels (Angelopoulos and Bates, 2021). Note that in the first stage, miscoverage refers to not sampling the acceptable responses, and this EER is controlled by $\alpha$. In the second stage, miscoverage refers to the calibrated prediction sets failing to cover the acceptable responses, and this EER is bounded by both $\alpha$ and $\beta$. Additionally, we adopt the commonly used Accuracy (ACC) and the Average Prediction Set Size (APSS) (Kostumov et al., 2024; Ye et al., 2024) to evaluate MLLM predictions. More details can be found in Appendix F.
@@ 4.2 Results for Risk Control
Following prior work (Mohri and Hashimoto, 2024; Su et al., 2024; Quach et al., 2024), we randomly split our datasets in half by default, calibrating the sampling number and threshold for identifying high-quality responses on the first half, i.e., the calibration set, and measure the EER (i.e., miscoverage rate) at various risk levels (i.e., upper bound) in two steps on the second half, i.e., the test set. We plot the empirical results in the two steps utilizing the GPT-4o mini model on four datasets in Figure 2(a) and Figure 2(b), and defer similar plots employing other MLLMs to Appendix G.
@@ 4.3 Results for Risk Assessment
Figure 3: Comparison of APSS before and after duplication at different risk levels ($\varepsilon$) across four open-source MLLMs on the VMME dataset.
@@
Conformal sets are constructed adaptively to each particular input (Angelopoulos and Bates, 2021), so that the set size gives us an indicator of the model’s uncertainty (or the risk in the current decision-making). Prior work (Ye et al., 2024; Kostumov et al., 2024) utilize APSS to benchmark LLMs and Vision-Language Models (VLMs) through risk assessment in MCQA tasks. In contrast to closed-ended scenarios, where MLLMs are provided a predetermined set of options that encompasses the correct answer and produce a prediction set without internal duplicates, open-ended VideoQA tasks feature their unfixed and unbounded output spaces. Consequently, duplicate responses are inevitable during the sampling process, which leads to semantic redundancy [^3] In the open-ended VideoQA task, sampling multiple responses from MLLMs can result in the candidate set containing identical answers or answers that are syntactically or lexically distinct but semantically equivalent. in the prediction set after filtering for high-quality responses based on the nonconformity score.
@@
Unlike prior work (Kostumov et al., 2024) that evaluates MLLMs on vision-language QA datasets, we investigate the impact of the audio modality on MLLM performance in open-ended VideoQA tasks. Specifically, we extract audio information from the original video, conduct varying degrees of Silence Percentage (SP), and then provide the processed audio along with visual and textual information to MLLMs. As illustrated in Figure 4, incorporating additional modality information enhances the ACC of MLLMs to some extent. As the SP increases, we observe significant changes in model uncertainty through the APSS metric, showcasing that introducing audio modality enhances the confidence level of MLLMs’ decision-making. In light of the observations mentioned above, we hope that our framework can assist in evaluating the performance of MLLMs in VideoQA tasks through risk assessment, while providing guarantees, thereby complementing the accuracy metric for a more comprehensive analysis.
@@ 4.4 Sensitivity Analyses
Conditional Miscoverage Rate.
As mentioned before, TRON bounds the correctness miscoverage rate marginally over the test set (i.e., on average). However, for some critical applications requiring more stringent guarantees, users demand that the error rate is bounded for a particular test data point. In other words, the risk level should be guaranteed for any subset of the data points. In the most general case, conditional coverage is impossible to achieve (Angelopoulos and Bates, 2021). Following prior work (Kumar et al., 2023; Su et al., 2024), we set $\alpha=\beta=0.1$ (i.e., $\varepsilon=0.19$) and evaluate the size-stratified miscoverage rate utilizing the Gemini-1.5-Pro model. As demonstrated in Figure 6, the miscoverage rate fluctuates up and down with the set size, with some values exceeding $\varepsilon$. It is worth noting that when we employ Semantic Diversity as the reliability measurement, TRON’s conditional performance improves (e.g., the max miscoverage rate decreases from 0.2065 to 0.1667 on the VMME dataset). At this point, it’s like a bag of tricks, and we can enhance the capability of reliability measurement to optimize the conditional performance of risk control.

### v3

@@ 1 Introduction
In recent years, there has been swift progress in the development of Large Language Models (LLMs) and Multimodal Large Language Models (MLLMs) (Zhao et al., 2023; Wu et al., 2023). MLLMs extend the capabilities of LLMs by integrating and processing information from multiple modalities, including text, vision, and audio (Bai et al., 2024; Zhu et al., 2024). However, MLLMs are proven to exhibit significant drawbacks in trustworthiness, such as hallucination (Rawte et al., 2023; Zhang et al., 2024; Wang et al., 2024a), which results in non-factual information (Ji et al., 2023) and biased generations (Zou et al., 2023). These issues have elicited increasing societal concerns regarding the reliable deployment of foundation models in consumer-facing applications (Qian et al., 2024).
Uncertainty estimation (Ye et al., 2024) provides valuable insights into the trustworthiness of model generations. Prior work performs probabilistic modeling (Li et al., 2024), develops entropy-based measures (Wang et al., 2024c; Duan et al., 2024; Farquhar et al., 2024), and generates multiple responses to analyze the output space (Lin et al., 2024; Wang et al., 2024b). However, these approaches cannot provide guarantees of the error rate, and forcing a measure to predict a single data point is overly restrictive (Cresswell et al., 2024). Split Conformal Prediction (SCP) (Papadopoulos et al., 2002; Angelopoulos & Bates, 2021; Campos et al., 2024) is a promising candidate to tackle these challenges above, which constructs statistically rigorous prediction sets with the hallmark of risk control (Angelopoulos et al., 2024; Farinhas et al., 2024). Furthermore, SCP allows the calibrated set size to vary based on the model’s uncertainty regarding a particular input.
SCP has been successfully applied in language modeling (Campos et al., 2024). However, existing methods either (1) modify the original model outputs to ensure factuality with desired probability (Mohri & Hashimoto, 2024; Cherian et al., 2024), potentially leading to uninformative and vague generations, (2) rely on internal token sequence logits (Quach et al., 2024), or (3) are restricted to multiple-choice settings (Kumar et al., 2023; Ye et al., 2024; Kostumov et al., 2024). Adapting SCP for proprietary MLLMs in practical open-ended video question-answering (VideoQA) applications presents several challenges: (a) Adaptability. Unlike closed-ended tasks equipped with fixed options covering the correct answers, the output space for each VideoQA task is unbounded, and there may not be an acceptable response within the sampled generations; (b) Reliability. Logits or verbalized confidence levels are often miscalibrated, leading to biased prediction sets; (c) Flexibility. For some API-only MLLMs, users lack access to internal model information like logits.
@@
Once the risk level for acceptable responses being sampled is specified, we evaluate the reliability of each response. Inspired by self-consistency theory (Wang et al., 2022; Li et al., 2022b), we propose using the frequency of each response as a proxy for confidence [^1] Response frequency positively correlates with average real generative probability. As the frequency increases, there is a corresponding growing level of confidence in the model’s responses (Su et al., 2024)., compensating for the limitations of internal logits information and verbalized confidence scores, and define the nonconformity score of each calibration data as one minus the frequency score of any one acceptable response in the candidate set. Then, we specify another risk level $\beta$ and calculate the approximate $1-\beta$ quantile of all nonconformity scores, denoted as $\hat{s}$. The quantile serves as a threshold to identify high-quality responses within the candidate set of each test data point. Finally, we guarantee that the miscoverage rate of acceptable responses in the calibrated prediction sets is statistically controlled by two risk levels, $\alpha$ and $\beta$ (i.e., $\leq\alpha+\beta-\alpha\beta$), while the set size reflects the model uncertainty to each query.
We evaluate our framework on two closed-ended VideoQA datasets (e.g., Video-MME (Fu et al., 2024) and NExT-QA (Xiao et al., 2021)), and two open-ended VideoQA datasets (e.g., MUSIC-AVQA (Li et al., 2022a) and MSVD (Chen & Dolan, 2011)), utilizing five open-source MLLMs(e.g., VideoLLaMA (Zhang et al., 2023), PandaGPT (Su et al., 2023), and NExT-GPT (Wu et al., 2024)), and three closed-source MLLMs(e.g., Gemini (Reid et al., 2024) and GPT-4o mini (OpenAI, 2023)). Experimental results demonstrate that TRON guarantees the error rates in both steps across various user-specified risk levels. Additionally, the deduplicated average set size provides stable uncertainty estimation of MLLMs across various risk levels in open-ended VideoQA tasks, complementing the accuracy metric for more comprehensive evaluations. Furthermore, we investigate other black-box measurements in the development of the nonconformity score, resulting in more efficient predictions and consistently rigorous guarantees. To summarize, our contributions are the following:
@@ 2 Related Work
Our work extends the variants of the SCP framework for correctness (coverage) guarantees and risk control in language modeling. Existing methods either modify model generations and achieve conformal factuality guarantees (Mohri & Hashimoto, 2024; Cherian et al., 2024) or guarantee the ground-truth coverage in Multiple-Choice Question-Answering (MCQA) tasks (Kumar et al., 2023; Ye et al., 2024; Kostumov et al., 2024). The previous approaches may result in vague and uninformative responses, constrained by the correctness of the original generation, while the latter frameworks are limited to closed-ended settings. In open-ended scenarios, (Quach et al., 2024) develops a stopping rule with slightly relaxed coverage criteria utilizing the internal logits information, restricting the applicability to API-only models, whereas (Su et al., 2024) leverages black-box uncertainty but fails to align the nonconformity score with correctness, leading to theoretically loose guarantees. Our approach addresses these gaps by introducing a novel two-step framework that first derives the minimum number of response samples to approximate a “closed-ended” condition, where there is at least one acceptable option provided, followed by identifying reliable responses based on self-consistency theory, which features (1) rigor, controlled by two user-specified risk levels, $\alpha$ and $\beta$, and (2) versatility, applicable to any MLLM that supports sampling in both closed-ended and open-ended settings.
@@
Recent work performs probabilistic modeling (Li et al., 2024), develops entropy-based uncertainty measures (Wang et al., 2024c; Duan et al., 2024; Farquhar et al., 2024), and generate multiple responses to analyze the output space (Lin et al., 2024; Wang et al., 2024b). However, these approaches are heuristic, and forcing a measure to predict a single data point is overly restrictive (Cresswell et al., 2024). (Ye et al., 2024) and (Kostumov et al., 2024) apply SCP to MCQA tasks, creating conformal sets based on a measure of uncertainty (risk), where the set size represents the confidence of the foundational model regarding a particular input, with larger sets indicating higher risk. In open-ended scenarios, there may be redundant responses within the prediction sets due to the randomness of sampling (Quach et al., 2024). Our work explores the performance of set size on risk assessment before and after deduplication and observes that the original set size becomes unstable in uncertainty estimation across different MLLMs as the error rate changes, while the calibrated set size reliably evaluates MLLMs across various risk levels.
@@ 3.1 Preliminaries
Following standard SCP for risk control (Angelopoulos & Bates, 2021; Angelopoulos et al., 2024), we assume a calibration dataset ${\mathcal{D}}_{cal}=\left\{\left(X_{i},Y_{i}\right)\right\}_{i=1}^{N}$ of $N$ query-answer pairs and an independent and identically distributed (i.i.d.) [^2] Split conformal prediction requires that the test data points are drawn from the same distribution as the calibration data points. test point $\left(x_{test},y_{test}\right)$. Our goal of risk control is to construct prediction sets that are calibrated to bound the correctness miscoverage rate
@@ 3.2 Conformal Risk Control
Inspired by self-consistency theory (Wang et al., 2022; Li et al., 2022b; Wang et al., 2024b), which states that a repetitively sampled response is viewed as a form of consistency linked to higher confidence in the response, we perform a simple semantic clustering process on the candidate set and obtain the frequency of each response, denoted as $F\left(\hat{y}_{m}^{(i)}\right)$. For a detailed description of semantic clustering, refer to Appendix C.1. Then, we define the nonconformity score of the $i$-th calibration data point as one minus the frequency of the acceptable response, denoted as $\hat{y}_{ref}^{(i)}$, within the candidate set, which is semantically equivalent to the label $Y_{i}$ (i.e., $\hat{y}_{ref}^{(i)}\Leftrightarrow Y_{i}$). $A\Leftrightarrow B$ represents a bidirectional entailment (MacCartney & Manning, 2015; Mohri & Hashimoto, 2024; Farquhar et al., 2024) between A and B (i.e., equivalence). The specific evaluation for bidirectional entailment is provided in Appendix C.1. In this case, the nonconformity score is formulated as
@@ 4.1 Experimental Settings
We employ five open-source MLLMs, including VideoLLaMA-7B (Zhang et al., 2023), VideoLLaMA-13B, PandaGPT-7B (Su et al., 2023), PandaGPT-13B, and NExT-GPT (Wu et al., 2024), and three closed-source MLLMs, including Gemini-1.5-Flash (Reid et al., 2024), Gemini-1.5-Pro, and GPT-4o mini (OpenAI, 2023). Note that we utilize all five open-source MLLMs as if they are API-only MLLMs, i.e., it assumes no access to any internal information of MLLMs.
@@
We consider two multiple-choice VideoQA datasets: Video-MME (VMME) (Fu et al., 2024) and NExT-QA (NEXT) (Xiao et al., 2021), where we only utilize the multiple-choice section of NEXT, and two open-ended VideoQA datasets: MUSIC-AVQA (MUSC) (Li et al., 2022a) and MSVD (Chen & Dolan, 2011). More details of the dataset utilization can be found in Appendix C.
@@
We utilize the Empirical Error Rate (EER) to assess whether we control the risk of miscoverage by the two user-specified risk levels (Angelopoulos & Bates, 2021). Note that in the first stage, miscoverage refers to not sampling the acceptable responses, and this EER is controlled by $\alpha$. In the second stage, miscoverage refers to the calibrated prediction sets failing to cover the acceptable responses, and this EER is bounded by both $\alpha$ and $\beta$. Additionally, we adopt the commonly used Accuracy (ACC) and the Average Prediction Set Size (APSS) (Kostumov et al., 2024; Ye et al., 2024) to evaluate MLLM predictions. More details can be found in Appendix F.
@@ 4.2 Results for Risk Control
Following prior work (Mohri & Hashimoto, 2024; Su et al., 2024; Quach et al., 2024), we randomly split our datasets in half by default, calibrating the sampling number and threshold for identifying high-quality responses on the first half, i.e., the calibration set, and measure the EER (i.e., miscoverage rate) at various risk levels (i.e., upper bound) in two steps on the second half, i.e., the test set. We plot the empirical results in the two steps utilizing the GPT-4o mini model on four datasets in Figure 2(a) and Figure 2(b), and defer similar plots employing other MLLMs to Appendix G.
@@ 4.3 Results for Risk Assessment
Figure 3: Comparison of APSS before and after deduplication at different risk levels ($\varepsilon$) across four open-source MLLMs on the VMME dataset.
@@
Conformal sets are constructed adaptively to each particular input (Angelopoulos & Bates, 2021), so that the set size gives us an indicator of the model’s uncertainty (or the risk in the current decision-making). Prior work (Ye et al., 2024; Kostumov et al., 2024) utilize APSS to benchmark LLMs and Vision-Language Models (VLMs) through risk assessment in MCQA tasks. In contrast to closed-ended scenarios, where MLLMs are provided a predetermined set of options that encompasses the correct answer and produce a prediction set without internal duplicates, open-ended VideoQA tasks feature their unfixed and unbounded output spaces. Consequently, duplicate responses are inevitable during the sampling process, which leads to semantic redundancy [^3] In the open-ended VideoQA task, sampling multiple responses from MLLMs can result in the candidate set containing identical answers or answers that are syntactically or lexically distinct but semantically equivalent. in the prediction set after filtering for high-quality responses based on the nonconformity score.
@@
Unlike prior work (Kostumov et al., 2024) that evaluates MLLMs on vision-language QA datasets, we investigate the impact of the audio modality on MLLM performance in open-ended VideoQA tasks. Specifically, we extract audio information from the original video, apply varying levels of Silence Percentage (SP), defined as the proportion of randomly muted audio segments in a video, and then provide the processed audio, along with the corresponding visual and textual information, to MLLMs. As illustrated in Figure 4, incorporating additional modality information enhances the ACC of MLLMs to some extent. As the SP increases, we observe significant changes in model uncertainty through the APSS metric, showcasing that introducing audio modality enhances the confidence level of MLLMs’ decision-making. In light of the observations mentioned above, we hope that our framework can assist in evaluating the performance of MLLMs in VideoQA tasks through risk assessment, while providing guarantees, thereby complementing the accuracy metric for a more comprehensive analysis.
@@ 4.4 Sensitivity Analyses
Conditional Performance.
As mentioned before, TRON bounds the correctness miscoverage rate marginally over the test set (i.e., on average). However, for some critical applications requiring more stringent guarantees, users demand that the error rate is bounded for a particular test data point. In other words, the risk level should be guaranteed for any subset of the data points. In the most general case, conditional coverage is impossible to achieve (Angelopoulos & Bates, 2021). Following prior work (Kumar et al., 2023; Su et al., 2024), we set $\alpha=\beta=0.1$ (i.e., $\varepsilon=0.19$) and evaluate the size-stratified miscoverage rate utilizing the Gemini-1.5-Pro model. As demonstrated in Figure 6, the miscoverage rate fluctuates up and down with the set size, with some values exceeding $\varepsilon$. It is worth noting that when we employ Semantic Diversity as the reliability measurement, TRON’s conditional performance improves (e.g., the max miscoverage rate decreases from 0.2065 to 0.1667 on the VMME dataset). At this point, it’s like a bag of tricks, and we can enhance the capability of reliability measurement to optimize the conditional performance of risk control.